OpenStudy (anonymous):

Angle A and B are supplementary, the measure of angle A is 2n and the measure of angle B is 3n + 2. Find the measure of angle B

OpenStudy (anonymous):

supplementary means sum has to be 180 deg so 2n + 3n + 2 = 180 5n = 180 - 2 n = 178/5 n = 35.6 deg